How much Na2O2 would need to react in order to release 1000 kJ of heat? Show your work.

2 Na2O2 (s) + 2 H2O (l) ---> 4NaOH (aq) + O2 (g) Delta Hrxn = -126 kJ
Answer #1

from the given reaction,

126 kJ of heat is released from 2 mole of Na2O2

therefore, to release 1000kJ of heat we need \frac{2}{126} \times 1000 = 15.87 mole of Na2O2

now molar mass of Na2O2 = 78 g

mass of 15.87 mole of Na2O2 = 15.87 * 78 = 1237.86 g of Na2O2

therefore, 1237.86 g of Na2O2 is required to produce 1000kJ of heat

note - delta Hrxn = negative means that heat is released.